# Hardware Co-Designed Optimal Control for Programmable Atomic Quantum Processors via Reinforcement Learning

### v1

@@ II.1 Formulating Co-designed Optimization Problem
We formalize the hardware co-designed QOC problem as a constrained optimization task, which aims to find the optimal control pulses applied to programmable units in the hardware to implement high-fidelity target gate operations. The entire workflow the co-designed QOC optimizer is illustrated in Fig. 1 (a), with details explained in the following. A mathematical model ($\mathscr{M}(\{\theta\},\{V(t)\})$ of the control hardware describes how the input laser fields $\{a_{in}\}$ are mapped to a set of dynamically modulated output fields $\{b_{out}(t)\}$ that address the involved atom qubits for gate implementation. Here, $\{\theta\}$ represent a set of design parameters characterizing the control hardware and $\{V(t)\}$ are the electrical signals applied to the programmable units in PIC for field modulation. Incorporating realistic hardware imperfections such as inter-channel crosstalk and beam leakage can be directly captured in this mathematical map. The output fields $\{b_{out}(t)\}$ further enter the control Hamiltonian $H_{control}(t)$ that determines the system evolution and final gate fidelity. The total system Hamiltonian is given as:
@@
static parameter $\text{static parameter sets}\quad\{\theta,\xi\}.$ (5)
@@ II.2 Physical System Model
Figure 2: (a) A 16-channel programmable atomic PIC control hardware fabricated using a 200 mm CMOS process [23]. Insets show SEM images of the piezo-actuated Si3N4 dual-ring Mach-Zehnder modulator (DRMZM). (b) Simulated inter-channel crosstalk arising from evanescent waveguide coupling in the PIC. Dotted curves represent data from 2D FDTD simulations using FlexCompute Tidy3D, assuming refractive indices of 2.0255 for Si3N4 and 1.45 for SiO2 at a wavelength of 780 nm. Coupling coefficient $\kappa_{0}=10.145$ and decay factor $\alpha=6.934$ are extracted from the fitted solid curve. (c, d) Heatmaps showing the amplitude (c) and phase (d) crosstalk matrices induced by waveguide coupling in a six-channel PIC. The matrices are generated using the fitted model from (b), assuming a coupling length of 600 $\mu$m and a channel pitch of 1.0 $\mu$m.
piezo-opto-mechanical actuation  [24]. Details of device structure and transfer function model of the DRMZM can be found in Ref. [21]. The inter-channel crosstalk effect induced by evanescent coupling between waveguides in channels $m$ and $n$ is captured using a coupling coefficient $C_{m,n}=A_{m,n}e^{i\theta_{m,n}}$. According to coupled mode theory  [25] under assumption of a symmetric directional coupler, we set amplitude $A_{m,n}$ as $\left|\sin\left(\kappa_{\text{eff},m,n}\,L_{m,n}\right)\right|$ and phase $\theta_{m,n}$ as $-\pi/2$. The effective coupling factor $\kappa_{\text{eff},m,n}$ is calculated as $\sqrt{\kappa_{m,n}^{2}+\Delta\beta^{2}}$, with the nominal coupling coefficient $\kappa_{m,n}=\kappa_{0}\exp\left(-\alpha\,d_{m,n}\right)$ and phase matching $\Delta\beta=\frac{\beta_{m}-\beta_{n}}{2}$. Here, the propagation constant $\beta_{i}$ for channel $i$ is given by $\frac{2\pi}{\lambda_{0}}\,n_{\text{eff},i}$ with $n_{eff,i}$ being the effective refractive index of the waveguide in channel $i$ and $\lambda_{0}$ being the wavelength of the propagating field , which is $780\,nm$ in our case. For $\kappa_{m,n}$, the prefactor $\kappa_{0}$ is the base coupling coefficient for nearest-neighboring waveguides, $\alpha$ is the decay factor that governs how quickly the coupling strength falls off with distance $d_{m,n}$ between channels $m$ and $n$, and $L_{m,n}$ is the coupling length between the involved waveguides. These two parameters, $\kappa_{0}$ and $\alpha$, can be obtained by fitting to 2D finite-difference time-domain (FDTD) simulation results obtained for neighboring $Si_{3}N_{4}$ waveguides, as seen in Fig. 2 (b). To account for practical fabrication variations of PICs in the foundry, the values of $d_{m,n}$ and $L_{m,n}$ are determined by combining a deterministic component with random variations. Specifically, the distance between channels is calculated as $d_{m,n}=|m-n|\times d_{0}+\delta_{d}$, where $d_{0}$ is the nominal channel pitch and $\delta_{d}$ is a small random variation uniformly sampled from a specified range. Similarly, the coupling length between channels is computed as $L_{m,n}=L_{0}\times s^{|m-n|}+\delta_{L}$, where $L_{0}$ is the nominal coupling length for waveguides in nearest-neighbor channels, $s$ is a scaling factor adjusting the nominal coupling length with increasing channel separation, and $\delta_{L}$ is a random perturbation uniformly sampled from a given range. As an example, Fig. 2(c) and (d) show the calculated amplitude and phase crosstalk matrices for a 6-channel photonic integrated circuit (PIC), where the nominal coupling length is $600\,\mu\text{m}$ (with $s$ set to 1.1 and $\delta_{L}$ uniformly sampled from $[-60\,\mu\text{m},\,60\,\mu\text{m}]$) and the channel spacing is $1.0\,\mu\text{m}$ (with $\delta_{d}$ uniformly sampled from $[-0.1\,\mu\text{m},\,0.1\,\mu\text{m}]$). Given the on-diagonal evolution represented by $T\{V(t)\}$, which corresponds to the model of channel-wise modulator unit, and the off-diagonal crosstalk coefficients $C_{m,n}$, the final unitary matrix modeling the PIC is constructed as:
@@ III.1 Hybrid SADE-Adam Approach
17:     Update: $x\leftarrow x-\alpha_{t}\,g$triangleright via Adam with clipping
@@ III.2 PPO-Based RL Approach
To formulate our hardware co-designed QOC problem as a RL problem, we cast the task of steering a quantum system toward a desired gate operation into the framework of PPO, which is a widely used policy-gradient algorithm designed to improve training stability while maintaining sample efficiency. Specifically, we design a custom Gymnasium environment, QOCEnv, where the state is defined by the historical control voltages and current fidelity metrics, and the action space comprises continuous voltage adjustments that are later scaled to the physical range of $[-15\,\mathrm{V},15\,\mathrm{V}]$. The reward function is carefully tailored to incentivize improvements in gate fidelity via adaptive scaling:
@@ IV Numerical Results and Discussions
We now present numerical results demonstrating the functionality of the hardware co-designed QOC framework. In Sec. IV.1, we first illustrate the necessity of employing QOC for mitigating hardware-related control imperfections by comparing traditional Gaussian-shaped pulses and optimal piecewise-shaped pulses obtained with the classical SADE-Adam optimizer, considering a simple program instruction with $N_{a}$=1. Then in Sec. IV.2, we benchmark the performance of three optimization methods, taking the classical hybrid SADE-Adam optimizer as a baseline and comparing conventional PPO-based RL optimizer and the End-to-End differentiable RL-based optimizer, assuming static control imperfections including inter-channel crosstalk and beam leakage based on models described in Sec. II. We then further validate the performance robustness of the End-to-End differentiable RL-based optimizer considering randomized dynamic control imperfections. Training for RL-based methods was conducted on a single NVIDIA A30 GPU, while the classical SADE-Adam optimizer was evaluated on a 16-core Intel Xeon CPU. In all simulations, following parameters are used: laser intensity of $20\ mW/cm^{2}$, detuning of $1\ GHz$, hyperfine splitting of $6.835\ GHz$, fixed gate duration of $0.1\mu s$, and a total of 100 numerical time steps.
@@ IV.1 Necessity of QOC
Figure 3: (a) Gaussian control pulse $V(t)$ applied only to channel 1 targeting atom 1, aiming to implement the gate $X_{1}/I_{2}/I_{3}$ on atoms 1, 2, and 3. (b) Optimized Gaussian control pulses $V(t)$ applied to all channels, mitigating beam leakage and inter-channel crosstalk to implement the same target gate as in (a). (c, d) Normalized amplitude (c) and phase (d) profiles of the optical field on the 2D atomic plane, taken at half the total gate operation time. Field profiles for cases (a), (b), and (c) are shown from left to right. The inter-atomic spacing is set to 3 $\mu$m, and the beam waist is 2 $\mu$m, resulting in significant leakage between atoms. (e) Histogram comparing gate fidelities achieved in cases (a), (b), and (c), highlighting the fidelity degradation due to beam leakage and crosstalk. (f) Gate error optimization progress curves for the same target gate $X_{1}/I_{2}/I_{3}$, using the classical hybrid SADE-Adam QOC optimizer. Each curve represents an independent test run initialized with a different random guess for the control pulses $V_{\text{ini}}(t)$. (g) Example of optimized control pulses $V(t)$ for channels 1–3 targeting atoms 1–3, from the first test run in (f). Control voltages are constrained within $[-15,+15]$ V, consistent with practical hardware limits.
@@ Appendix B PPO-based RL Approach
To apply PPO to the QOC problem, we construct a custom Gymnasium environment called QOCEnv, where the agent optimizes the control signal set $\{V(t)\}$ that programs the light fields applied to atoms to implement the gate operation and thereby determine the gate fidelity. The state consists of the history of control voltages over past time steps along with the current fidelity progress. The action space comprises continuous voltage adjustments in the range $[-1,1]$, which are later scaled to the physical range of $[-15\,\mathrm{V},15\,\mathrm{V}]$. The reward function is carefully tailored to incentivize fidelity improvement via adaptive scaling:
@@
The PPO agent is trained over 10,000 episodes with a learning rate of $1\times 10^{-4}$, an entropy coefficient of $0.1$ to encourage exploration, a discount factor of $\gamma=0.99$, a clipping threshold of $0.2$, generalized advantage estimation (GAE) with $\lambda=0.95$, and a maximum gradient norm of $0.5$ for gradient clipping. During training, the RL-based optimizer iteratively refines the control signals to maximize gate fidelity. An adaptive logging system tracks fidelity improvements and saves the corresponding voltage sequences whenever a new maximum fidelity is achieved. Training terminates when either the fidelity reaches $F\geq 0.999$ or if the best fidelity stagnates for over 1000 episodes within the range $0.9990.99\leq F<$ 0.999. Additionally, the adaptive reward scaling mechanism can detect prolonged stagnation and trigger early termination.
@@ Appendix C End-to-End Differentiable RL Approach
The policy network is a minimal yet expressive multilayer perceptron (MLP) designed to generate correlated and structured control sequences across all channels. It consists of two fully connected layers with 64 neurons each, using tanh activation functions to enforce smooth control variations. The input to the network is a latent noise vector $z$, sampled from a standard normal distribution, which enables the model to explore a diverse set of control strategies. The network transforms this latent input into structured control voltages using a hierarchical mapping. The first layer applies a non-linear transformation to the latent vector, capturing global correlations in the control sequence. The second layer refines these representations, ensuring that the output follows physically meaningful constraints. The final output layer produces an $N_{\text{ch}}\times N_{\text{seg}}$ matrix, where each element represents the control voltage for a specific channel at a given time segment. The output values are linearly scaled to match the predefined voltage range $[-15V,15V]$, fulfilling the hardware imposed constraints. This formulation allows the policy network to learn temporally and spatially correlated control sequences, optimizing the overall quantum evolution trajectory.
@@ Acknowledgments
Qian Ding acknowledge funding support from the Swiss National Science Foundation (SNSF) Postdoc Mobility Fellowship. Dirk Englund acknowledge funding from NSF CQN. We also acknowledge the support from subMIT cluster at MIT for offering access to high-performance computational resources.

### v2

@@ II.1 Formulating Co-designed Optimization Problem
We formalize the hardware co-designed QOC problem as a constrained optimization task, which aims to find the optimal control pulses applied to programmable units in the hardware to implement high-fidelity target gate operations. The entire workflow the co-designed QOC optimizer is illustrated in Fig. 1 (a), with details explained in the following. A mathematical model $\mathscr{M}(\{\theta\},\{V(t)\})$ of the control hardware describes how the input laser fields $\{a_{in}\}$ are mapped to a set of dynamically modulated output fields $\{b_{out}(t)\}$ that address the involved atom qubits for gate implementation. Here, $\{\theta\}$ represent a set of design parameters characterizing the control hardware $\{V(t)\}$ are the electrical signals applied to the programmable units in PIC for field modulation. Incorporating realistic hardware imperfections such as inter-channel crosstalk and beam leakage can be directly captured in this mathematical map. The output fields $\{b_{out}(t)\}$ further enter the control Hamiltonian $H_{control}(t)$ that determines the system evolution and final gate fidelity. The total system Hamiltonian is given as:
@@
$\text{static parameter sets}\quad\{\theta,\xi\}.$ (5)
@@ II.2 Physical System Model
[…]
@@ III.1 Hybrid SADE-Adam Approach
17:  Update: $x\leftarrow x-\alpha_{t}\,g$ $\triangleright$ via Adam with clipping
@@ III.2 PPO-Based RL Approach
To formulate our hardware co-designed QOC problem as a RL problem, we cast the task of steering a quantum system toward a desired gate operation into the framework of PPO, which is a widely used policy-gradient algorithm designed to improve training stability while maintaining sample efficiency. Specifically, we design a custom Gymnasium environment, QOCEnv, where the state is defined by the historical control voltages and current fidelity metrics, and the action space comprises continuous voltage adjustments that are later scaled to the physical range of $15\,\mathrm{V},15\,\mathrm{V}]$. The reward function is carefully tailored to incentivize improvements in gate fidelity via adaptive scaling:
@@ IV Numerical Results and Discussions
We now present numerical results demonstrating the functionality of the hardware co-designed QOC framework. In Sec. IV.1, we first illustrate the necessity of employing QOC for mitigating hardware-related control imperfections by comparing traditional Gaussian-shaped pulses and optimal piecewise-shaped pulses obtained with the classical SADE-Adam optimizer, considering a simple program instruction with $N_{a}$=1. Then in Sec. IV.2, we benchmark the performance of three optimization methods, taking the classical hybrid SADE-Adam optimizer as a baseline and comparing conventional PPO-based RL optimizer and the End-to-End differentiable RL-based optimizer, assuming static control imperfections including inter-channel crosstalk and beam leakage based on models described in Sec. II. We then further validate the performance robustness of the End-to-End differentiable RL-based optimizer considering randomized dynamic control imperfections. Training for RL-based methods was conducted on a single NVIDIA A30 GPU, while the classical SADE-Adam optimizer was evaluated on a 16-core Intel Xeon CPU. In all simulations, following parameters are used: laser intensity of $20\ mW/cm^{2}$, detuning of 1GHz1\ GHz, hyperfine splitting of 6.835GHz6.835\ GHz, fixed gate duration of $0.1\mu s$, and a total of 100 numerical time steps.
@@ IV.1 Necessity of QOC
an example, we consider a easy task, where the program instruction is to apply X-gate only on atom 1, and we can construct $U_{\text{target}}=X_{1}\otimes I_{2}\otimes I_{3}$. In an ideal scenario without inter-channel crosstalk or beam leakage, applying a Gaussian $\pi$-pulse exclusively to atom 1 while keeping the control signals for atoms 2 and 3 at zero, as shown in Fig. 3 (a), would implement $U_{\text{target}}$ with high fidelity, as the control field would not influence neighboring atoms. However, when atoms arranged in a triangular subgroup with a spacing of 3 µm and a beam waist of 2 µm, the control field applied to atom 1 inevitably leaks onto atoms 2 and 3, as illustrated in the left plots of the field’s amplitude and phase profiles in Fig. 3 (c) and (d). In this case, noted as case (a), applying a Gaussian control pulse solely to atom 1 , results in a significantly low fidelity of $\sim 0.262$, as seen in the first histogram bar in Fig. 3 (e). To mitigate this undesired beam leakage effect, compensatory control pulses must be applied to atoms 2 and 3 to preserve their identity gates. In this case, noted as case (b), with optimized Gaussian control signals shown in Fig. 3 (b) applied to all atoms, the fidelity improves considerably to $\sim 0.952$, as indicated by the second histogram bar in Fig. 3 (e). The corresponding field profiles in this case are given by the middle plots in Fig. 3 (c) and (d), where both amplitude and phase distributions change significantly compared with case (a). However, when further including inter-channel crosstalk due to waveguide coupling, additional distortions of control fields profiles are induced, see the right plots in Fig. 3 (c) and (d), compared with middle plots for case (b). In this case, noted as case (c), the gate fidelity drops dramatically down to $\sim 0.059$, as shown in third histogram bar in Fig. 3(e). While additional fine-tuning of control pulses for atoms 2 and 3 could partially mitigate these additional field distortions, the complexity of the interference patterns induced by the inter-channel crosstalk grows as the control system scales up, where simple reshaping of Gaussian control pulses becomes very limited to counteract intricate crosstalk effects in a systematic way. Instead, these challenges can be effectively addressed using the hardware co-designed QOC approach, which optimizes the set of control signals $\{V(t)\}$ to implement the desired $U_{\text{target}}$. To demonstrate this, we present results obtained using the classical hybrid SADE-Adam optimizer, which iteratively refines the control signals to maximize fidelity. The inter-channel crosstalk and beam leakage are considered the same as in the case (c) using Gaussian pulses. The progress curves of fidelity optimization in Fig. 3 (f) show rapid fidelity improvements with all final fidelities above 99.9% (errors below $1\text{e-}3$ shown in the plot) across multiple trials with different initial guesses of control signals $\{V_{\text{ini}}(t)\}$, highlighting the effectiveness and robustness of this approach. The final optimized control signals from one representative trial (Set 1) are shown in Fig. 3 (g). The piecewise-shaped control signals applied to DRMZMs in channels 0–2 manipulate the fields addressing atoms 1–3 and $V_{0,1}$ in each channel represent signals applied to the two rings in a DRMZM of the PIC.
@@ Appendix B PPO-based RL Approach
To apply PPO to the QOC problem, we construct a custom Gymnasium environment called QOCEnv, where the agent optimizes the control signal set $\{V(t)\}$ that programs the light fields applied to atoms to implement the gate operation and thereby determine the gate fidelity. The state consists of the history of control voltages over past time steps along with the current fidelity progress. The action space comprises continuous voltage adjustments in the range [1,1][-1,1], which are later scaled to the physical range of $15\,\mathrm{V},15\,\mathrm{V}]$. The reward function is carefully tailored to incentivize fidelity improvement via adaptive scaling:
@@
The PPO agent is trained over 10,000 episodes with a learning rate of $1\times 10^{-4}$, an entropy coefficient of $0.1$ to encourage exploration, a discount factor of $\gamma=0.99$, a clipping threshold of $0.2$, generalized advantage estimation (GAE) with $\lambda=0.95$, and a maximum gradient norm of $0.5$ for gradient clipping. During training, the RL-based optimizer iteratively refines the control signals to maximize gate fidelity. An adaptive logging system tracks fidelity improvements and saves the corresponding voltage sequences whenever a new maximum fidelity is achieved. Training terminates when either the fidelity reaches $F\geq 0.999$ or if the best fidelity stagnates for over 1000 episodes within the range $0.99\leq F<0.999$. Additionally, the adaptive reward scaling mechanism can detect prolonged stagnation and trigger early termination.
@@ Appendix C End-to-End Differentiable RL Approach
The policy network is a minimal yet expressive multilayer perceptron (MLP) designed to generate correlated and structured control sequences across all channels. It consists of two fully connected layers with 64 neurons each, using tanh activation functions to enforce smooth control variations. The input to the network is a latent noise vector $z$, sampled from a standard normal distribution, which enables the model to explore a diverse set of control strategies. The network transforms this latent input into structured control voltages using a hierarchical mapping. The first layer applies a non-linear transformation to the latent vector, capturing global correlations in the control sequence. The second layer refines these representations, ensuring that the output follows physically meaningful constraints. The final output layer produces an $N_{\text{ch}}\times N_{\text{seg}}$ matrix, where each element represents the control voltage for a specific channel at a given time segment. The output values are linearly scaled to match the predefined voltage range [15V,15V][-15V,15V], fulfilling the hardware imposed constraints. This formulation allows the policy network to learn temporally and spatially correlated control sequences, optimizing the overall quantum evolution trajectory.
@@ Acknowledgments
Qian Ding acknowledge funding support from the Swiss National Science Foundation (SNSF) Postdoc Mobility Fellowship. Dirk Englund acknowledge funding from NSF CQN. We acknowledge the support from the NVIDIA Academic Grant and subMIT cluster at MIT for offering access to high-performance computational resources.